AGw Commissioning with IRTC 02-03 November 2008 MST

Observers: D.Thompson at LBT, J.Storm from Potsdam
Software Support: S.Hooper, N. Cushing C.Biddick on call as needed.
Telescope Operator: D. Gonzalez Huerta

Summary

Details (times reported in UT below)

Weather at sunset, some cirrus low on S horizons, otherwise clear. 5.5C, 10-12m/s wind, H=60% and rising.

Z4/Z11 sign test for Andrew:

UT01:15
PSFL M1 Z(total)=2.133, add 1000nm Z4 and Z(total)=2.178mm
PSFL M2 Z(total)=0.000, add 1000nm Z4 and Z(total)=-0.038mm

PSFL M1 Z(total)=2.132, add 1000nm Z11 and Z(total)=2.223
PSFL M2 Z(total)=0.000, add 1000nm Z11 and Z(total)=0.000

Correct pointing, initial collimation

IE=-43.0 CA=-31.0 at start.
01:26 pointto WT10_304 gs=0 Tel="ACQUIRE" PAM-
  • Initial preset failed: Nov 2 18:25:30 lbtmu105 LBT_IIF: Invalid buffer: RPCERROR
  • Resend same preset succeeded at 18:33:57

M1 Zglobal adjusted to 0.32mm
IE=-34.0 CA=-25.0 (corrected at EL~82, Tout=5.1C)

Check pointing at a lower elevation:
01:43 pointto WT10_236 gs=0 Tel="ACQUIRE" PAM-
IE=-39.0 CA=-29 (EL~64, Tout=5.1C) This is 4 arcsec IE and 2 arcsec CA off from last night.

pointto BS9183 gs=0 Tel="ACTIVE"
GCS is seeing ~1.7 arcsec FWHM, ~250nm rms per WFS cycle

UT 02:26 Humidity still climbing, 70%, Tout=4.8, dewpoint depression is 5C.

Recheck AGW rotation center

Repeat end of last night measurements:
  • collimate: Lpoint M39collim cat="AGW_collim.cat" WLT=0.63
  • stop guiding/wfs
  • setxy -x 0.0 -y 612.5
  • take 4+ sets of 270 degree rotations, 60s exposure

02:42 Data in /Repository/20081103_IRTC/
guider000001 is 5s TRACK mode on the M39 field
guider000002 60s while rotating from 217 -> -55, Center at 288 249
guider000003 60s while rotating from -55 -> 217, Center at 288 248
guider000004 60s while rotating from 217 -> -55, Center at 287 248
guider000005 60s while rotating from -55 -> 217, Center at (can't measure)
guider000006 60s while rotating from 217 -> -55, Center at 287 248
guider000007 60s while rotating from -55 -> 217, Center at 287 248
Note: Seeing is poor during this sequence (hard to eyeball center of star trails). The center of rotation is consistent across the 5 usable measurements (287,248) and within ~4 pixels of the nominal AGW hotspot location (289,252).

Preset back to collimate here failed (star too faint). Source in acquisition image, but strongly comatic. Why did it change by so much?

Humidity 80% and still climbing.

pointto BS9188 gs=0 - preset successful. Also heavily comatic. First WFS=1943nm, second=1344nm, third=728nm

03:10 We have 84% humidity and rising rapidly, 70% heavy cirrus coverage, and poor seeing. IR Satellite shows more coming, so we are closing for the time being and will monitor conditions.

05:34 91% humidity and still rising, dewpoint depression of 1.3C. Needless to say we are still closed.

Jesper checked in, but we are still closed.

10:30 100% humidity. Stayed that way the rest of the night, darn it.

No luck getting anything below here:

  • IRTC orientation to N
  • If both OK, can start mapping AGW field transformation
  • Proper pointing model data?

-- DavidThompson - 02 Nov 2008
Topic revision: r7 - 05 Nov 2008, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback